OpenStudy (anonymous):

solve the following equations using the method indicated? a. substitution y=3x+4 y=2x+6

OpenStudy (anonymous):

first one? What's the question?

OpenStudy (anonymous):

what do you mean

OpenStudy (anonymous):

the first part of the question doesn't make sense... is it y=3x=4 or y=3x-4?

OpenStudy (anonymous):

+ sorry ...

OpenStudy (anonymous):

Ok y=3x+4 y=2x+6 2x+6=3x+4 (subsitute y in second equation into first) 2x-3x=4-6 -x=-2 x=2

OpenStudy (anonymous):

and y is

OpenStudy (anonymous):

y= 3(2) +4 y=6+4 y=10 (subsitute x value in any of two equations)

OpenStudy (anonymous):

i really don't get any of this... i suck at math period...

OpenStudy (anonymous):

because both equations are equivalent (equal one another) You can replace the first equation into the second... and the y value in the second equation can replace the y value in the first equation... then you have two equaions which you can solve for y... then u get the y value and can alternatively choose either equation to substitute that value to get the x valuel...

OpenStudy (anonymous):

sorry i don't get get any of this

OpenStudy (anonymous):

awwww. mmmm it's substitution... I had trouble with it at first. Try asking a teacher for more help or anyone here on basic rules for substitution. I'm tired and am off to bed.. Sorry I can;t help you more :(

OpenStudy (anonymous):

the thing is my exam is tomorrow :(